OpenStudy (anonymous):

somebody gave me an answer earlier but i dont get it im sorry Your classmate solved (x+5)(x-1)=2 as follows: x+5=2 x-1=2 x=2-5 x=2+1 x= -3 x=3 Your teacher said this was wrong because the values did not check. Explain what is wrong with your classmate’s solution.

OpenStudy (anonymous):

Thats wrong

OpenStudy (anonymous):

whats wrong?

OpenStudy (somy):

x+5=2 x-1=2 if these were to get you 2 then in the answer u'd end up with 4

OpenStudy (somy):

and your answer has to be 2 not 4

hero (hero):

For one... \((-3 + 5)(-3 - 1) \ne 2\)

hero (hero):

\((3 + 5)(3 - 1) \ne 2\) either

OpenStudy (jack1):

(x+5)(x-1)=2 so use FOIL to solve (x+5)(x-1)=2 x^2 + 5x - 1x - 5 = 2 x^2 + 4x - 7 = 0 now use the quadratic equation to solve for x you should arrive at x = 1.3 ish and -5.3 ish

OpenStudy (anonymous):

In order to solve such problems we ned to multiply both the brackets first and then bring 2 to left side. (x+5)(x-1)=2 x(x-1) +5(x-1) -2=0 \[\huge x^2 -x + 5x- 5 -2 =0\] \[\huge x^2 + 4x- 7=0\] now factorize it and you will find the correct answer,

OpenStudy (anonymous):

thank you guys so much :)

hero (hero):

Whenever you have an equation of the form (x - m)(x - n) = p You essentially have a quadratic equation since the product of two binomials will always form a quadratic expression. Furthermore after expanding the product of two binomials, you must put it in form \(ax^2 + bx + c = 0\) before attempting to solve it. In other words, the left side must be a quadratic expression and the right side must be equal to zero BEFORE attempting to solve for x. If you proceed in any other manner, you will get the wrong answer.

OpenStudy (anonymous):

okay thank you :)

hero (hero):

Also you should know that if what you had was of the form \((x - m)(x - n) = 0\), then you would be able to use the zero product property to split the equation in to two equations such as x - m = 0 x - n = 0

hero (hero):

Again, you can only do this when the product of two binomials equal zero.

OpenStudy (anonymous):

yes i know that thanks :)

hero (hero):

If that's the case, and you know what to do now, how would you go about solving (x - 3)(x + 4) = 5 ?

OpenStudy (anonymous):

its just like the problem below

hero (hero):

I'd like to see you attempt to solve the problem of course just to make sure you know what you are doing.

OpenStudy (anonymous):

im going to finish this one first okay (the problem below) im not yet done with it :(

hero (hero):

What problem are you referring to?

OpenStudy (anonymous):

the one i posted

hero (hero):

Post it here.

OpenStudy (anonymous):

okaaay

OpenStudy (anonymous):

\[\frac{ x=-4+\sqrt{4^{2}}-(4)(1)(-7) }{ 2 (1) }\]

OpenStudy (anonymous):

and \[\frac{ x=-4 - \sqrt{4^{2}}-(4)(1)(-7) }{ 2(1) }\]

OpenStudy (anonymous):

eq1: \[\frac{ x=-4+\sqrt{44} }{ 2(1) }\]

OpenStudy (anonymous):

eq2: \[\frac{ x=-4-\sqrt{44} }{ 2(1) }\]

OpenStudy (anonymous):

eq1: 1.32 & eq2: -5.32 (rounded off)

OpenStudy (anonymous):

am i correct or not :c

hero (hero):

Well, I think you meant to put it in this form: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Perhaps you should review a few lessons o how to post fractions and equations using \(\LaTeX\)

OpenStudy (anonymous):

ohmygod what is dat omg im only a 9th grader :(((((

OpenStudy (anonymous):

but is my answer correct?